# Boundary-value problems of functional differential equations with state-dependent delays

### v2

@@ 2.3 Illustrative example of BVP
We will use the example (1.3) throughout to illustrate concepts and results. The embedded BVP for finding periodic orbits of this FDE, rescaled to base interval $[0,1]$ with a possible choice for additional affine conditions is
$y^{\prime}(t)$ $=-Ty(t-(p+y(t))/T),\mbox{\quad($y(t)=y(t+1)$ for all $t\in\mathbb{R}$),}$, (2.5)
00 0 $=R_{\mathrm{aff}}[y,T,p]:=\begin{bmatrix}y(0)\\ 2\int_{0}^{1}\sin(2\pi t)y(t)\mathop{}\!\mathrm{d}t-y_{0}\end{bmatrix}$ (2.6)
for a range of small $y_{0}$. It is sufficient to impose (2.5) on the base interval $[0,1)$, if one takes into account the periodicity of $y$. For the choice (2.6) of $R_{\mathrm{aff}}$ the first condition fixes the phase of the solution $y$ (serving as a phase condition) and the second condition fixes the amplitude, locally parametrizing the solution family by the constant $y_{0}$ (thus, serving as a pseudo-arclength condition). System (2.5)–(2.6) will have solutions of the form $y(t)=y_{0}\sin(2\pi t)+O(|y_{0}|^{2})$, $T=2\pi+O(|y_{0}|^{2})$, $p=\pi/2+O(|y_{0}|^{2})$ for small $y_{0}$ according to the Hopf bifurcation theorem Sieber (2012). The unknowns in this problem are $(y(\cdot),T,p)$.
@@ 3 Fixed-point problem equivalent to periodic BVP and relevant notation
$y^{\prime}(t)&=G(y_{t},\mu),&&\in\mathbb{R}^{n_{y}}\mbox{\ for $t\in[0,1)$,}&&\mbox{(FDE for $1$-periodic continuous $y$),}\\ 0&=R_{\mathrm{aff}}[y,\mu]&&\in\mathbb{R}^{n_{\mu}}&&\mbox{(affine constraints).}\end{aligned}$ (3.2)
@@
[…]
$C^{k,1}_{\pi}$ =$=\{v\in C^{k}_{\pi},\operatorname{Lip}v^{(k)}<\infty\},$ $\|y\|_{k,1}$ $\!=\max\{\|v\|_{k},\operatorname{Lip}v^{(k)}\},$
$L^{\infty}_{\pi}$ =$=\{v\mbox{\ ess.bd.,\ }v(t)=v(t+1)\mbox{\ for all $t$}\},$ $\|v\|_{\infty}$ $=\operatorname{ess\,sup}_{t\in[0,1]}|v(t)|.$
@@
$g$ $:C^{0,1}_{\mathrm{e}}\to C^{0}_{\mathrm{e}}\mbox{,\quad with}$ $g\left(\begin{bmatrix}v\\ \alpha\\ \mu\end{bmatrix}\right)$ $=\begin{bmatrix}t\mapsto G(v_{t},\mu)\\ \alpha\\ \mu+R_{\mathrm{aff}}[v,\mu]\end{bmatrix}\mbox{,\quad and}$ (3.6)
$\mathcal{L}$ :$&L^{\infty}_{\mathrm{e}}\to C^{0,1}_{\mathrm{e}}\\ \makebox[0.0pt]{or\qquad}&C^{\ell}_{\mathrm{e}}\to C^{\ell+1}_{\mathrm{e}}\end{aligned}\mbox{,\quad with}$ $\mathcal{L}\begin{bmatrix}w\\ \alpha\\ \nu\end{bmatrix}\hskip 8.00003pt$ $=\begin{bmatrix}\displaystyle t\mapsto\alpha+\int_{0}^{t}w(s)\mathop{}\!\mathrm{d}s-t\int_{0}^{1}w(s)\mathop{}\!\mathrm{d}s\\[5.69054pt] \alpha+\displaystyle\int_{0}^{1}w(s)\mathop{}\!\mathrm{d}s\\ \nu\end{bmatrix}.$ (3.7)
@@
For functions $z\in C^{0}_{\pi}$ we define the interpolation projection $P_{L}z$ as the unique piecewise polynomial on mesh $(t_{i})_{i=0}^{L}$ of degree $m-1$ where the piece on each interval $(t_{i-1},t_{i})$ equals $z$ on the nodes $t_{i,j}$:
@@
$P_{L}z$ $=\hat{z}\quad\begin{aligned} &\mbox{with $\hat{z}(t_{i,j})=z(t_{i,j})$ for all $i\in\{1,\ldots,L\},j\in\{1,\ldots,m\}$,}\\ &\mbox{and\ $\hat{z}$ is degree $m-1$ polynomial on $(t_{i-1},t_{i})$ for all $i\in\{1,\ldots,L\}$.}\end{aligned}$
In the name $P_{L}$ we do not indicate the dependence on the interpolation degree $m-1$ as we will keep this degree constant, studying only the limit $L\to\infty$ in our convergence analysis. The dependence on the mesh $(t_{i})$, which will change with increasing $L$, is also implicitly included in the subscript $L$. The interpolating piecewise polynomial $t\mapsto[P_{L}z](t)$ is not necessarily continuous as it may have discontinuities at the mesh boundaries $t_{i}$, such that the codomain of $P_{L}$ is $L^{\infty}_{\pi}$. By construction of $\Phi=\mathcal{L}\circ g$ and $P_{L}$ we have the following equivalence.
@@ 4 Mildly differentiable nonlinear functionals
The regularity properties outlined in this section imply, in particular, that we cannot rely on the general framework by Maset Maset (2016) to prove the convergence of the numerical method, in contrast to Andò and Breda (2020), which studied the constant delay case. In other words, it is not possible to prove all the (theoretical and numerical) assumptions made in Maset (2016) to reach our convergence result. In particular, some assumptions will only hold in a weaker form, with different norms in the relevant inequalities than those in the general framework in Maset (2016). In section 6 we will show that this does not impede high-order convergence for polynomial collocation methods, and eventually get to convergence results identical to those for constant delays.
@@ 5 Assumptions on the (infinite-dimensional) problem and immediate consequences
This section reformulates Assumption 2.2, stated in Section 2 for the BVP, as assumptions on the infinite-dimensional fixed-point problem (3.5) for $\Phi$. These are fewer than stated for the general theory by Maset Maset (2016), because Maset includes further assumptions needed for determining the convergence rate of the Newton iterations when solving the discretized problem.
@@
such that $v^{*}$ will even be in $C^{1}_{\pi}$. The Hopf bifurcation theorem, proved for functional differential equations with state-dependent equations in (Sieber, 2012), provides a scenario that ensures the existence of periodic orbits, and, hence, solutions of fixed-point problem (3.5). The illustrative example (1.3) is chosen to satisfy the assumptions of the Hopf bifurcation theorem for $p\approx\pi/2$ and small-amplitude periodic solutions $y(t)$ with period $2\pi$ (hence, solutions $(y(\cdot),T,p)$ of (2.5), (2.6) with small-amplitude $y$).
@@
The right-hand side $G$ of the DDE (3.2) is mildly differentiable to order ${\ell_{\max}}\geq 1$.
@@
Let $x^{*}=(v^{*},\alpha^{*},T^{*})$ be a solution of the fixed-point problem (3.5), and let the right-hand side $G$ of the DDE (3.2) satisfy mild differentiability to order ${\ell_{\max}}$. Then the solution component $v^{*}$ is in $C^{{\ell_{\max}}+1}_{\pi}$ as a function of time, and, hence, $x^{*}\in C^{{\ell_{\max}}+1}_{\mathrm{e}}$.
The statement of Corollary 5.3 follows from Lemma 4.1, applied to the case $\ell=0$, and using that $v^{*}$ satisfies the differential equation (5.1): for each $j$ from 00 to ${\ell_{\max}}$ we have that $G(v^{*}_{t},\mu^{*})$ is in $C^{j}_{\pi}$ because $v^{*}$ is in $C^{j}_{\pi}$ by Lemma 4.1. Then by the differential equation (5.1) $(v^{*})^{\prime}(t)$ is in $C^{j}_{\pi}$, such that $v^{*}$ is in $C^{j+1}_{\pi}$.
Hence, by Assumption 5.2 that ${\ell_{\max}}\geq 1$, we have that $x^{*}\in C^{{\ell_{\max}}+1}_{\mathrm{e}}\subseteq C^{2}_{\mathrm{e}}$. Note that Assumption 5.2 of mild differentiability of $G$ with ${\ell_{\max}}=1$, together with Corollary 5.3, imply that the Fréchet derivative
@@
Elements $y$ of the nullspace of $[I-D\Phi(x^{*})]$ that are at least in $C^{0}_{\mathrm{e}}$ satisfy the identity $y=\mathcal{L}Dg(x^{*})y$, such that they are in the image of $\mathcal{L}$ of $C^{0}_{\mathrm{e}}$, which is in $C^{1}_{\mathrm{e}}$, since $\mathcal{L}$ involves an anti-derivative. Hence, we may replace $C^{1}_{\mathrm{e}}$ in Assumption 5.4 by $C^{0}_{\mathrm{e}}$ (and, even $L^{\infty}_{\mathrm{e}}$).
Since $D\Phi(x^{*})$ is a compact linear operator on spaces $C^{k,j}_{\mathrm{e}}$ for all $k\geq 0$ and $j=0,1$, the nullspace of $D\Phi(x^{*})$ is at most finite-dimensional, and $\dim\ker[I-D\Phi(x^{*})]=0$ implies the existence of a bounded inverse $[I-D\Phi(x^{*})]^{-1}$. A sequence of results formulate Assumption 5.4 in terms of requiring full rank of a finite-dimensional characteristic matrix. One example construction for a characteristic matrix of a linear time-periodic FDE is given by Sieber and Szalai (2011), generalized to FDEs of neutral type by Verduyn Lunel (2023). The construction of $\Delta(\lambda)$ in (Sieber and Szalai, 2011) is primarily useful for proving the existence of a characteristic matrix for periodic problems. In practice numerical methods for determining eigenvalues of monodromy operators, such as the pseudospectral method (Breda et al., 2006, 2012), work directly on large matrices generated by the discretization projection $\mathcal{P}_{L}$, without prior reduction. Borgioli et al. (2020) proves convergence of methods as implemented in DDE-Biftool and knut.
@@ 6 Convergence of solutions of discretized problem
In this section, we consider that solutions $x$ of the fixed point problem for the discretized map given in the Equivalence Lemma 3.1,
@@
Furthermore, the interpolation approximation $\mathcal{P}_{L}$ satisfies for any continuous function $z$ (see (Rivlin, 1969))
@@
Under Assumptions 5.1, 5.2, 5.4 from section 5, the fixed point problem $\Phi_{L}(x)=x$ is stable in $x^{*}$, that is, there exists a bound $C_{\mathrm{stab},L}$ for $\|[I-D\Phi_{L}(x^{*})]^{-1}\|_{C^{0,1}_{\mathrm{e}}\leftarrow C^{0,1}_{\mathrm{e}}}$ for sufficiently large $L$ such that $C_{\mathrm{stab},L}\to C_{\mathrm{stab},\infty}$ for $L\to\infty$.
@@
By Assumption 5.4, $I-D\Phi(x^{*})$ has a bounded inverse. The Banach perturbation lemma (e.g., (Ortega, 1990, Theorem 2.1.1)) and Lemma 6.1 then ensure that $I-D\Phi_{L}(x^{*})=I-\mathcal{L}\mathcal{P}_{L}Dg(x^{*})$ has a bounded inverse, too, for sufficiently large $L$. Let $L_{\mathrm{diff}}$ be such that $\omega^{*}(1/L_{\mathrm{diff}})\leq\/(2C_{\mathrm{stab},\infty})$. Then, for $L\geq L_{\mathrm{diff}}$, the inverse has the norm
@@
where $m-1$ is the degree of the interpolation polynomial used in $\mathcal{P}_{L}$ in (3.9).
@@
By Corollary 5.3, $x^{*}$ is in $C^{{\ell_{\max}}+1}_{\mathrm{e}}$ and, thus, $g(x^{*})$ is in $C^{\ell_{\max}}_{\mathrm{e}}$. Let $[t_{i},t_{i+1}]$ be one of the subintervals in the mesh of the discretization, where $i\in\{0,\ldots,L-1\}$. If ${\ell_{\max}}\geq m$, then
@@
$\|(\mathcal{P}_{L}-I)g(x^{*})|_{[t_{i},t_{i+1}]}\|_{\infty}\leq c\left[\frac{t_{i+1}-t_{i}}{2m}\right]^{-{\ell_{\max}}}\|g(x^{*})^{({\ell_{\max}})}\|_{0}.$
@@
[…]
$\|\epsilon_{\mathrm{nl}}(L,\delta^{x})\|_{0,1}$ $\leq\rho_{\mathrm{nl}}\|\delta^{x}\|_{0,1}\mbox{\quad for all $\delta^{x}\in B^{0,1}_{r_{\mathrm{nl}}}(0)$.}$
@@
for all $\delta^{x,1},\delta^{x,2}\in B^{0,1}_{r_{\mathrm{nl}}}(0)\subset C^{0,1}_{\mathrm{e}}$. In the denominator on the right-hand side the norm $\|\mathcal{P}_{L}\|_{L^{\infty}_{\mathrm{e}}\leftarrow C^{0}_{\mathrm{e}}}$ has a uniform upper bound for all $L$ as we keep the degree $m-1$ of the interpolation polynomial in $\mathcal{P}_{L}$ in (3.9) fixed. Note also the stronger $\|\cdot\|_{0,1}$-norm on the right-hand side in (6.13) required by mild differentiability. Thus, by definition of $\epsilon_{\mathrm{nl}}$ in (6.4)
@@
Let $G$ be mildly differentiable to order ${\ell_{\max}}\geq 1$, and let $x^{*}=(v^{*},\alpha^{*},\mu^{*})$ be a fixed point of $\Phi$, as defined in (3.5), with bounded inverse of $I-D\Phi(x^{*})$, as listed in Assumptions  5.1, 5.2, 5.4 in Section 5.
@@
$L_{\min}=$ $\max\{L_{\mathrm{diff}},L_{\mathrm{c}}\}\mbox{,\quad where $L_{\mathrm{c}}$ is s.\,t.\quad}\|\epsilon_{\mathrm{c}}(L)\|_{0,1}\leq\cfrac{1-\kappa}{C_{\mathrm{stab}}}\,r_{\mathrm{nl}}\mbox{\ for all $L\geq L_{\mathrm{c}}$.}$
@@
Revisiting the definition (6.10) for the bound $C_{\mathrm{stab}}$ in Corollary 6.3, we observe that we can replace $C_{\mathrm{stab}}$ by the smaller $C_{\mathrm{stab},L}$, which can be chosen as close to the infinite-dimensional stability constant $C_{\mathrm{stab},\infty}$ as one wishes if one increases the discretization level $L$ further. Similarly, we may choose the constant $\kappa$ as close to 00 as desired, where again a smaller $\kappa$ requires a larger lower bound on $L$. Furthermore, assuming that the order of mild differentiability ${\ell_{\max}}$ exceeds the order $m$ of the discretization, we may insert the concrete estimate (6.12) for $\epsilon_{\mathrm{c}}(L)$. We also observe that the $m$th time derivatives $[g(x^{*})]^{(m)}$ equal the derivatives of order $m+1$ of $v^{*}$, when we denote the first component of the solution $x^{*}$ as $v^{*}$ ($x^{*}=(v^{*},\alpha^{*},\mu^{*})$). With these concrete estimates we obtain that for every $\epsilon>0$ there exists a lower bound $L_{\mathrm{low}}(\epsilon)$ such that the error of the discretized fixed point problem $\delta^{x,L}$ satisfies

### v3

@@ 2.3 Illustrative example of BVP
We will use the example (1.3) throughout to illustrate concepts and results. The embedded BVP for finding periodic orbits of (1.3), rescaled to base interval $[0,1]$ with additional affine conditions is
$y^{\prime}(t)$ $=-Ty(t-(p+y(t))/T),\mbox{\quad($y(t)=y(t+1)$ for all $t\in\mathbb{R}$),}$ (2.4)
$0$ $=R_{\mathrm{aff}}[y,T,p]:=\begin{bmatrix}y(0)\\ 2\int_{0}^{1}\sin(2\pi t)y(t)\mathop{}\!\mathrm{d}t-y_{0}\end{bmatrix}$ (2.5)
for a range of small $y_{0}$. It is sufficient to impose (2.4) on the base interval $[0,1)$, if one takes into account the periodicity of $y$. For the choice (2.5) of $R_{\mathrm{aff}}$ the first condition fixes the phase of the solution $y$ (serving as a phase condition) and the second condition fixes the amplitude, locally parametrizing the solution family by the constant $y_{0}$ (thus, serving as a pseudo-arclength condition). System (2.4)–(2.5) will have solutions of the form $y(t)=y_{0}\sin(2\pi t)+O(|y_{0}|^{2})$, $T=2\pi+O(|y_{0}|^{2})$, $p=\pi/2+O(|y_{0}|^{2})$ for small $y_{0}$ according to the Hopf bifurcation theorem (Sieber, 2012). The unknowns in this problem are $(y(\cdot),T,p)$.
@@ 3 Fixed-point problem equivalent to periodic BVP and relevant notation
$\begin{aligned} y^{\prime}(t)&=G(y_{t},\mu),&&\in\mathbb{R}^{n_{y}}\mbox{\ for $t\in[0,1)$,}&&\mbox{(FDE for $1$-periodic continuous $y$),}\\ 0&=R_{\mathrm{aff}}[y,\mu]&&\in\mathbb{R}^{n_{\mu}}&&\mbox{(affine constraints).}\end{aligned}$ (3.2)
@@
$C^{k\phantom{,1}}_{\pi}$ $\!=\{v:\mbox{\ $k\times$\,cont.\,diff.,\,}v(t)=v(t+1)\mbox{\ for all $t$}\},$ $\|v\|_{k\phantom{,1}}$ $\!=\max_{t\in[0,1],j\leq k}|v^{(j)}(t)|,$
$C^{k,1}_{\pi}$ $=\{v\in C^{k}_{\pi},\operatorname{Lip}v^{(k)}<\infty\},$ $\|y\|_{k,1}$ $\!=\max\{\|v\|_{k},\operatorname{Lip}v^{(k)}\},$
$L^{\infty}_{\pi}$ $=\{v\mbox{\ ess.bd.,\ }v(t)=v(t+1)\mbox{\ for all $t$}\},$ $\|v\|_{\infty}\$$=\operatorname{ess\,sup}_{t\in[0,1]}|v(t)|.$
@@
[…]
$\mathcal{L}$ $:\begin{aligned} &L^{\infty}_{\mathrm{e}}\to C^{0,1}_{\mathrm{e}}\\ \makebox[0.0pt]{or\qquad}&C^{\ell}_{\mathrm{e}}\to C^{\ell+1}_{\mathrm{e}}\end{aligned}\mbox{,\quad with}$ $\mathcal{L}\begin{bmatrix}w\\ \alpha\\ \nu\end{bmatrix}\hskip 8.00003pt$ $=\begin{bmatrix}\displaystyle t\mapsto\alpha+\int_{0}^{t}w(s)\mathop{}\!\mathrm{d}s-t\int_{0}^{1}w(s)\mathop{}\!\mathrm{d}s\\[5.69054pt] \alpha+\displaystyle\int_{0}^{1}w(s)\mathop{}\!\mathrm{d}s\\ \nu\end{bmatrix}$. (3.7)
@@
For functions $z\in C^{0}_{\pi}$ we define the interpolation projection $P_{L}z$ as the unique piecewise polynomial on mesh $(t_{i})_{i=0}^{L}$ of degree m1m-1 where the piece on each interval $(t_{i-1},t_{i})$ equals $z$ on the nodes $t_{i,j}$:
@@
[…]
In the name $P_{L}$ we do not indicate the dependence on the interpolation degree m1m-1 as we will keep this degree constant, studying only the limit $L\to\infty$ in our convergence analysis. The dependence on the mesh $(t_{i})$, which will change with increasing $L$, is also implicitly included in the subscript $L$. The interpolating piecewise polynomial $t\mapsto[P_{L}z](t)$ is not necessarily continuous as it may have discontinuities at the mesh boundaries $t_{i}$, such that the codomain of $P_{L}$ is $L^{\infty}_{\pi}$. By construction of $\Phi=\mathcal{L}\circ g$ and $P_{L}$ we have the following equivalence.
@@ 4 Mildly differentiable nonlinear functionals
The regularity properties outlined in this section imply, in particular, that we cannot rely on the general framework by Maset (2016) to prove the convergence of the numerical method, in contrast to (Andò and Breda, 2020), which studied the constant delay case. In other words, it is not possible to prove all the (theoretical and numerical) assumptions made in (Maset, 2016) to reach our convergence result. In particular, some assumptions will only hold in a weaker form, with different norms in the relevant inequalities than those in the general framework in (Maset, 2016). In section 6 we will show that this does not impede high-order convergence for polynomial collocation methods, such that we obtain convergence results identical to those for constant delays.
@@ 5 Assumptions on the (infinite-dimensional) problem and immediate consequences
This section reformulates Assumption 2.2, stated in Section 2 for the BVP, as assumptions on the infinite-dimensional fixed-point problem (3.5) for $\Phi$. These are fewer than stated for the general theory by Maset (2016), because Maset includes further assumptions needed for determining the convergence rate of the Newton iterations when solving the discretized problem.
@@
such that $v^{*}$ will even be in $C^{1}_{\pi}$. The Hopf bifurcation theorem, proved for functional differential equations with state-dependent equations in Sieber (2012), provides a scenario that ensures the existence of periodic orbits, and, hence, solutions of fixed-point problem (3.5). The illustrative example (1.3) is chosen to satisfy the assumptions of the Hopf bifurcation theorem for $p\approx\pi/2$ and small-amplitude periodic solutions $y(t)$ with period $2\pi$ (hence, solutions $(y(\cdot),T,p)$ of (2.4), (2.5) with small-amplitude $y$).
@@
The right-hand side $G$ of the FDE (3.2) is mildly differentiable to order ${\ell_{\max}}\geq 1$.
@@
Let $x^{*}=(v^{*},\alpha^{*},T^{*})$ be a solution of the fixed-point problem (3.5), and let the right-hand side $G$ of the FDE (3.2) satisfy mild differentiability to order $\ell_{\max}$}. Then the solution component $v^{*}$ is in $C^{{\ell_{\max}}+1}_{\pi}$ as a function of time, and, hence, $x^{*}\in C^{{\ell_{\max}}+1}_{\mathrm{e}}$.
The statement of Corollary 5.3 follows from Lemma 4.1, applied to the case $\ell=0$, and using that $v^{*}$ satisfies the differential equation (5.1): for each $j$ from 0 to $\ell_{\max}$} we have that $G(v^{*}_{t},\mu^{*})$ is in $C^{j}_{\pi}$ because $v^{*}$ is in $C^{j}_{\pi}$ by Lemma 4.1. Then by the differential equation (5.1) $(v^{*})^{\prime}(t)$ is in $C^{j}_{\pi}$, such that $v^{*}$ is in $C^{j+1}_{\pi}$.
Hence, by Assumption 5.2 that ${\ell_{\max}}\geq 1$, we have that $x^{*}\in C^{{\ell_{\max}}+1}_{\mathrm{e}}\subseteq C^{2}_{\mathrm{e}}$. Note that Assumption 5.2 of mild differentiability of $G$ with ${\ell_{\max}}=1$, together with Corollary 5.3, imply that the derivative
@@
Elements $y$ of the nullspace of $I-D\Phi(x^{*})]$ that are at least in $C^{0}_{\mathrm{e}}$ satisfy the identity $y=\mathcal{L}Dg(x^{*})y$, such that they are in the image of $\mathcal{L}$ of $C^{0}_{\mathrm{e}}$, which is in $C^{1}_{\mathrm{e}}$, since $\mathcal{L}$ involves an anti-derivative. Hence, we may replace $C^{1}_{\mathrm{e}}$ in Assumption 5.4 by $C^{0}_{\mathrm{e}}$ (and, even $L^{\infty}_{\mathrm{e}}$).
Since $D\Phi(x^{*})$ is a compact linear operator on spaces $C^{k,j}_{\mathrm{e}}$ for all $k\geq 0$ and $j=0,1$, the nullspace of $D\Phi(x^{*})$ is at most finite-dimensional, and $\dim\ker[I-D\Phi(x^{*})]=0$ implies the existence of a bounded inverse $I-D\Phi(x^{*})]^{-1}$. A sequence of results formulate Assumption 5.4 in terms of requiring full rank of a finite-dimensional characteristic matrix. One example construction for a characteristic matrix of a linear time-periodic FDE is given by Sieber and Szalai (2011), generalized to FDEs of neutral type by Verduyn Lunel (2023). In practice numerical methods for determining eigenvalues of monodromy operators, such as the pseudospectral method (Breda et al., 2022, 2006, 2012), work directly on large matrices generated by the discretization projection $\mathcal{P}_{L}$, without prior reduction. Borgioli et al. (2020) proves convergence of methods as implemented in DDE-Biftool and knut.
@@ 6 Convergence of solutions of discretized problem
In this section, we show that solutions $x$ of the fixed point problem for the discretized map given in the Equivalence Lemma 3.1,
@@
Furthermore, the interpolation approximation $\mathcal{P}_{L}$ satisfies for any continuous function $z$ (see Rivlin (1969))
@@
Under Assumptions 5.1, 5.2, 5.4 from section 5 the fixed point problem $\Phi_{L}(x)=x$ is stable in $x^{*}$, that is, there exists a bound $C_{\mathrm{stab},L}$ for $\|[I-D\Phi_{L}(x^{*})]^{-1}\|_{C^{0,1}_{\mathrm{e}}\leftarrow C^{0,1}_{\mathrm{e}}}$ for sufficiently large $L$, and $C_{\mathrm{stab},L}\to C_{\mathrm{stab},\infty}$ for $L\to\infty$.
@@
By Assumption 5.4, $I-D\Phi(x^{*})$ has a bounded inverse. The Banach perturbation lemma (e.g., (Ortega, 1990, Theorem 2.1.1)) and Lemma 6.1 then ensure that $I-D\Phi_{L}(x^{*})=I-\mathcal{L}\mathcal{P}_{L}Dg(x^{*})$ has a bounded inverse, too, for sufficiently large $L$. Let $L_{\mathrm{diff}}$ be such that $\omega^{*}(1/L_{\mathrm{diff}})\leq 1/(2C_{\mathrm{stab},\infty})$. Then, for $L\geq L_{\mathrm{diff}}$, the inverse has the norm
@@
where m1m-1 is the degree of the interpolation polynomial used in $\mathcal{P}_{L}$ in (3.9).
@@
By Corollary 5.3, $x^{*}$ is in $C^{{\ell_{\max}}+1}_{\mathrm{e}}$ and, thus, $g(x^{*})$ is in $C^{\ell_{\max}}_{\mathrm{e}}$. Let $t_{i},t_{i+1}]$ be one of the subintervals in the mesh of the discretization, where $i\in\{0,\ldots,L-1\}$. If ${\ell_{\max}}\geq m$, then
@@
$\|(\mathcal{P}_{L}-I)g(x^{*})|_{[t_{i},t_{i+1}]}\|_{\infty}\leq c\left[\frac{t_{i+1}-t_{i}}{2m}\right]^{{\ell_{\max}}}\|g(x^{*})^{({\ell_{\max}})}\|_{0}.$
@@
$\|\epsilon_{\mathrm{nl}}(L,\delta^{x,1})-\epsilon_{\mathrm{nl}}(L,\delta^{x,2})\|_{0,1}$ $\leq\rho_{\mathrm{nl}}\|\delta^{x,1}-\delta^{x,2}\|_{0,1}\mbox{, implying, in particular,}$
$\|\epsilon_{\mathrm{nl}}(L,\delta^{x})\|_{0,1}$ $\leq\rho_{\mathrm{nl}}\|\delta^{x}\|_{0,1}\mbox{\qquad for all $\delta^{x}\in B^{0,1}_{r_{\mathrm{nl}}}(0)$.}$
@@
for all $\delta^{x,1},\delta^{x,2}\in B^{0,1}_{r_{\mathrm{nl}}}(0)$. In the denominator on the right-hand side the norm $\|\mathcal{P}_{L}\|_{L^{\infty}_{\mathrm{e}}\leftarrow C^{0}_{\mathrm{e}}}$ has a uniform upper bound for all $L$ as the degree m1m-1 of the interpolation polynomial in $\mathcal{P}_{L}$ in (3.9) is fixed. Note the stronger $\|\cdot\|_{0,1}$-norm on the right-hand side in (6.13) required by mild differentiability. Thus, by definition of $\epsilon_{\mathrm{nl}}$ in (6.4)
@@
Let $G$ be mildly differentiable to order ${\ell_{\max}}\geq 1$, and let $x^{*}=(v^{*},\alpha^{*},\mu^{*})$ be a fixed point of $\Phi$, as defined in (3.5), with bounded inverse of $I-D\Phi(x^{*})$, as listed in Assumptions 5.1, 5.2, 5.4 in Section 5.
@@
[…]
@@
Revisiting the definition (6.10) for the bound $C_{\mathrm{stab}}$ in Corollary 6.3, we observe that we can replace $C_{\mathrm{stab}}$ by the smaller $C_{\mathrm{stab},L}$, which can be chosen as close to the infinite-dimensional stability constant $C_{\mathrm{stab},\infty}$ as one wishes if one increases the discretization level $L$ further. Similarly, we may choose the constant $\kappa$ as close to 0 as desired, where again a smaller $\kappa$ requires a larger lower bound on $L$. Furthermore, assuming that the order of mild differentiability $\ell_{\max}$} exceeds the order $m$ of the discretization, we may insert the concrete estimate (6.12) for $\epsilon_{\mathrm{c}}(L)$. We also observe that the $m$th time derivatives $g(x^{*})]^{(m)}$ equal the derivatives of order $m+1$ of $v^{*}$, when we denote the first component of the solution $x^{*}$ as $v^{*}$ ($x^{*}=(v^{*},\alpha^{*},\mu^{*})$). With these concrete estimates we obtain that for every $\epsilon>0$ there exists a lower bound $L_{\mathrm{low}}(\epsilon)$ such that the error of the discretized fixed point problem $\delta^{x,L}$ satisfies